OpenStudy (anonymous):

I'm not very good with decimals. can someone help me solve by factoring? (0.3x-0.1)(0.05x+1)=0 I don't want just the answer but if it can be explained step by step please

OpenStudy (callisto):

Should I start from here :(0.3x-0.1)(0.05x+1)=0 ?

OpenStudy (anonymous):

I think im suppose to solve for both X's

OpenStudy (callisto):

Okay, the first thing you need to know is that no matter what times 0, it gives 0 as a result. Got this part?

OpenStudy (anonymous):

the first x = one third, the second x = negative twenty

OpenStudy (anonymous):

yes

OpenStudy (callisto):

So, you've got 2 factors now. (0.3x-0.1) and (0.05x+1). Multiplying both of them, it gives 0. So, either one of them is 0. Understand?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Then, put it as a equation Either (0.3x-0.1) = 0 or (0.05x+1)=0 Understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do i get rid of the decimals?

OpenStudy (callisto):

Now, you can solve the equations respectively. (0.3x-0.1) = 0 0.3 x =0.1 <- understand this step?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats about as far as i got and i did similar with the other equation too but im stuck at the decimals

OpenStudy (callisto):

Now, you need you know that 0.1 = 1/10 ... Can you understand this?

OpenStudy (anonymous):

ok

OpenStudy (callisto):

and you need to know that 0.1 x10 =1. Okay?

OpenStudy (anonymous):

ok

OpenStudy (callisto):

So 0.3 x 10 =?

OpenStudy (anonymous):

so i can just do to one side what i do to the other . thus multiplying ten on each side equaling one third

OpenStudy (anonymous):

thank you i think i got it, i gave you a medal

OpenStudy (callisto):

Wow~~ That's the concept! For 0.3 x =0.1 Multiplying both sides by 10, you'll get 3x =1 You can solve it :)

OpenStudy (callisto):

Welcome :)